You can use a minus sign to make a negative number like -2. What happens for each of the following and why?


>>> 2++2


>>> 2--2


>>> 2+-2


>>> 2-+2

1
Expert's answer
2022-04-15T04:43:29-0400

2++2=2+2=4

2—2=2+2=4

2+-2=2-2=0

2-+2=2-2=0